One sample has n=6 scores with a mean of . A second sample has n=4 scores with a mean of  . If the two samples are combined, what is the mean for the combined sample?


1
Expert's answer
2021-01-13T14:25:05-0500

In given question mean of samples are not given.

so let us assume One sample has n=6 scores with a mean of 12 . A second sample has n=4 scores with a mean of 16 . If the two samples are combined, what is the mean for the combined sample?


Formula: mean of combined sample = "\frac{m*X_a+n*X_b}{m+n}"


Where:

  • xa = the mean of the first set,
  • m = the number of items in the first set,
  • xb = the mean of the second set,
  • n = the number of items in the second set,
  • xc the combined mean.

Solution:

Mean of combined sample= "\frac{6*12+4*16}{6+4}"

="\frac{136}{10}=1.36"